OpenStudy (anonymous):

5x-4y=-18?? i need help ... ill give a medal

OpenStudy (anonymous):

are you trying to find y or x.?

OpenStudy (anonymous):

both its a linear equation

hero (hero):

Where's the other equation?

OpenStudy (anonymous):

there isnt one i have to come up with 3 solutions for x and y and then plot it on a graph

hero (hero):

Start with x = 0, x = 1, x = 2 Input those values into the equation to find the y values of each.

OpenStudy (anonymous):

okaii soo 5*1-4y=-18 do i divide the 5*1 ??

hero (hero):

No, you isolate y

OpenStudy (anonymous):

..... ummm

hero (hero):

I kept making sign mistakes

hero (hero):

The correct expression for y is \[y = \frac{5x + 18}{4}\]

hero (hero):

That's about five do-overs for me.

OpenStudy (anonymous):

my computer won't let me post the way i found answer

OpenStudy (anonymous):

but you solve 5x - 4y = -18 for x. You get x = (4/5)y - (18/5). Substitute 0 in for y. x= -3.6. substitute -3.6 in for x in the beginning equation. y = 0

hero (hero):

If you use x = 2, x = 6, x = 10, then you should get even values for y after inputting into \[y = \frac{5x + 18}{4}\]